# Uniform convergence of the smooth calibration error and its relationship with functional gradient

### v2

@@ C.3 Proof of Lemma 2
Similar to the proof of Lemma 1 in Appendix C.2, we use McDiarmid’s inequality to evaluate the exponential moment. For this purpose, we need to bound the sensitivity constants in Eq. (115). Recall the definition:
@@
$\phi(f,Z_{i},Z^{\prime}_{i})\coloneqq\left[(y_{i}-f(x_{i}))\left(\frac{1+\sigma_{i}}{2}\omega_{i}+\frac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\frac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\frac{1-\sigma_{i}}{2}\omega_{i}\right)\right].$ (139)
Note that in the special case where $\sigma_{i}=1$ for all $i\in[n]$, this quantity reduces to the difference of smooth CEs:
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\phi(f,Z_{i},Z^{\prime}_{i})=\lambda\left(\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\right).$ (140)
We now fix the dataset and consider changing a single coordinate $\sigma_{j}$ from $1$ to $-1$. This change affects only the $j$-th term of the sum. Specifically, for $\sigma_{j}=1$, the $j$-th term is
@@
and for $\sigma^{\prime}_{j}=-1$, the term becomes
@@
This corresponds to a swap of the training and test inputs in the smooth CE optimization.
To formalize this, define the datasets:
@@
Then we can write the exponent, where $\sigma_{i}=1$ for $i\neq j$ and $\sigma_{j}=-1$, as
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i\neq j}\phi(f,Z_{i},Z^{\prime}_{i})-\frac{\lambda}{n}\phi(f,Z_{j},Z^{\prime}_{j})=\lambda\left(\mathrm{smCE}(f,\tilde{S}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})\right).$ (141)
Using Eq. (130), which bounds the stability of $\mathrm{smCE}$, and combining with Eq. (140), we have:
@@
$=\lambda|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,\tilde{S}_{n})|+\lambda|\mathrm{smCE}(f,S^{\prime}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})|$ (143)
$\leq\frac{2\lambda}{n}.$ (144)
@@
The same argument applies to any fixed realization $\sigma=\sigma|_{\pm}$ of the Rademacher variables ((Here, $\sigma|_{\pm}$ is the realization, so it is not a random variable)). For each such realization, we can construct the corresponding datasets $S_{\sigma|_{\pm}}$ and $S^{\prime}_{\sigma|_{\pm}}$ used in the smooth CE optimization. Then:
@@
and changing any single component of $\sigma$ results in a difference bounded by $\frac{2\lambda}{n}$, as in Eq. (145).
Therefore, following the argument in Appendix C.2, we obtain that the sensitivity coefficients satisfy $c_{i}=\frac{2\lambda}{n}$ for all $i\in[n]$.
Substituting this into Lemma 4 in Appendix C.2 yields the desired exponential moment bound.

### v4

@@ C.3 Proof of Lemma 2
Similar to the proof of Lemma 1 in Appendix C.2, we apply McDiarmid’s inequality to evaluate the exponential moment. To this end, we first compute the bounded difference constants in Eq. equation 117. Recall the definition:
@@
$\phi(f,Z_{i},Z^{\prime}_{i})\coloneqq\left[(y_{i}-f(x_{i}))\left(\tfrac{1+\sigma_{i}}{2}\omega_{i}+\tfrac{1-\sigma_{i}}{2}\omega^{\prime}_{i}\right)-(y^{\prime}_{i}-f(x^{\prime}_{i}))\left(\tfrac{1+\sigma_{i}}{2}\omega^{\prime}_{i}+\tfrac{1-\sigma_{i}}{2}\omega_{i}\right)\right].$ (141)
For clarity, we first analyze the special case where $\sigma_{i}=1$ for all $i\in[n]$; the general case is handled later. Under this assumption, the expression reduces to the difference of smooth CEs:
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i=1}^{n}\phi(f,Z_{i},Z^{\prime}_{i})=\lambda\big(\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,S^{\prime}_{n})\big).$ (142)
Now fix the dataset and flip a single coordinate $\sigma_{j}$ from $1$ to 1-1. Only the $j$-th term is affected. For $\sigma_{j}=1$, the term is
@@
while for $\sigma_{j}=-1$, it becomes
@@
This corresponds to swapping the training and test inputs in the smooth CE optimization.
Formally, define the datasets:
@@
Then the exponent, with $\sigma_{i}=1$ for $i\neq j$ and $\sigma_{j}=-1$, can be written as
$\max_{\omega}\min_{\omega^{\prime}}\frac{\lambda}{n}\sum_{i\neq j}\phi(f,Z_{i},Z^{\prime}_{i})-\frac{\lambda}{n}\phi(f,Z_{j},Z^{\prime}_{j})=\lambda\big(\mathrm{smCE}(f,\tilde{S}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})\big).$ (143)
Using Eq. equation 132, which bounds the stability of $\mathrm{smCE}$, and combining with Eq. equation 142, we obtain:
@@
$=\lambda\big|\mathrm{smCE}(f,S_{n})-\mathrm{smCE}(f,\tilde{S}_{n})\big|+\lambda\big|\mathrm{smCE}(f,S^{\prime}_{n})-\mathrm{smCE}(f,\tilde{S}^{\prime}_{n})\big|$ (145)
$\leq\tfrac{2\lambda}{n}.$ (146)
@@
The same reasoning applies to any fixed realization $\sigma=\sigma|_{\pm}$ of the Rademacher variables (note that $\sigma|_{\pm}$ is a fixed realization and not a random variable). For each such realization, we construct the datasets $S_{\sigma|_{\pm}}$ and $S^{\prime}_{\sigma|_{\pm}}$ used in the smooth CE optimization. Then,
@@
and changing any single component of $\sigma$ alters the value by at most $\frac{2\lambda}{n}$, as in Eq. equation 147.
Therefore, by the argument in Appendix C.2, the sensitivity coefficients satisfy $c_{i}=\frac{2\lambda}{n}$ for all $i\in[n]$. Substituting this into Lemma 4 in Appendix C.2 yields the desired exponential moment bound.
@@ C.4 Proof of Corollary 1
Proof.
They further showed that there exists a universal constant $C$ such that
based on the result of ambroladze2004complexity. See also luxburg2004distance for the derivation of Rademacher complexity of lipshictz functions.
We now allocate the total failure probability $\delta$ across the two terms using the union bound. Specifically, we set $\delta\to\frac{2}{3}\delta$ in Eq. equation 149, and $\delta\to\frac{1}{3}\delta$ in Theorem 1 or Corollary 6. Then, the sum of the confidence terms becomes:
$\sqrt{\frac{\log\frac{2}{\frac{2}{3}\delta}}{2n}}+2\sqrt{\frac{\log\frac{1}{\frac{1}{3}\delta}}{n}}=\left(2+\frac{1}{\sqrt{2}}\right)\sqrt{\frac{\log\frac{3}{\delta}}{n}}\leq 3\sqrt{\frac{\log\frac{3}{\delta}}{n}}.$ (152)
Thus, we obtain the desired high-probability bound. ∎
Appendix D Analysis based on Rademacher complexity
D.1 Proof of Theorem 2
To simplify the notation, we define
$\phi(f,h,x,y)=h(f(x))\cdot(y-f(x)).$ (153)
Then
$\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})$ (154)
$\leq\sup_{f\in\mathcal{F}}\mathrm{smCE}(f,\mathcal{D})-\mathrm{smCE}(f,S_{\mathrm{tr}})$ (155)
$\leq\sup_{f\in\mathcal{F}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\mathbb{E}\phi(f,h,X,Y)-\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (156)
$\leq\sup_{f\in\mathcal{F}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\left\{\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\right\}$ (157)
$\leq\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\left\{\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\right\}.$ (158)
In the last line, we simply used the swapping of the supremum, which can be shown using the definition of the supremum. (In general, for any $f\in\mathcal{F}$, $\phi(f,h)\leq\sup_{f}^{\prime}\phi(f^{\prime},h)$, which leads to $\sup_{h}\phi(f,h)\leq\sup_{h}\sup_{f}^{\prime}\phi(f^{\prime},h)$. This further leads to $\sup_{f}\sup_{h}\phi(f,h)\leq\sup_{h}\sup_{f}^{\prime}\phi(f^{\prime},h)$.)
Then, we use Mcdiramid’s inequality. The bounded difference for the training dataset is $2/n$, we have the following result; for any $\delta>0$ with probability $1-\delta/2$, we have
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (159)
$\leq\mathbb{E}_{S_{\mathrm{tr}}\sim D^{n}}\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (160)
Then by considering the standard symmetrization property, we have the following upper bound
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})$ (161)
$\leq 2\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (162)
Finally, by considering the union bound, for any $\delta>0$ with probability $1-\delta$, we have
$\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\Big|\mathbb{E}\phi(f,h,X,Y)-\frac{1}{n}\sum_{i=1}^{n}\phi(f,h,X_{i},Y_{i})\Big|$ (163)
$\leq 2\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]+2\sqrt{\frac{\log\frac{2}{\delta}}{n}}.$ (164)
Next, we evaluate the Rademacher complexity term.
For the notational convenience, we define
$T$ $:=\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right],$ (165)
$\eta_{0}$ $:=\{\eta:[0,1]\!\to\!\mathbb{R}\ \mid\ -1\leq\mathrm{Lip}(\eta)\leq 1\}.$
Our goal is to show
$T\leq 2{\mathfrak{R}}_{\mathcal{D},n}(\mathcal{F})+\frac{12}{\sqrt{n}}\int_{0}^{1}\sqrt{\log N\!\bigl(u,\eta_{0},\|\cdot\|_{\infty}\bigr)}du.$ (166)
First, we will consider an $\epsilon$-net of the Lipschitz class, which is independent of data and $\sigma$. Fix geometric scales $\varepsilon_{k}:=2^{-k}$ for $k=0,1,2,\dots$. For each $k$, choose a fixed (data- and $\sigma$-independent) $\varepsilon_{k}$-net $\mathcal{N}_{k}\subset\eta_{0}$ under the uniform norm $\|\cdot\|_{\infty}$, with $|\mathcal{N}_{k}|=N(\varepsilon_{k},\eta_{0},\|\cdot\|_{\infty})$. For every $\eta\in\eta_{0}$ pick a nearest projection $\pi_{k}(\eta)\in\mathcal{N}_{k}$ so that $\|\eta-\pi_{k}(\eta)\|_{\infty}\leq\varepsilon_{k}$. Then, by the triangle inequality,
$\bigl\|\pi_{k}(\eta)-\pi_{k-1}(\eta)\bigr\|_{\infty}\leq\varepsilon_{k}+\varepsilon_{k-1}\leq 2\varepsilon_{k-1}.$ (167)
Consequently,
$\eta=\pi_{0}(\eta)+\sum_{k\geq 1}\bigl(\pi_{k}(\eta)-\pi_{k-1}(\eta)\bigr).$ (168)
We set $A(h)=\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})$. The discretization error by the $k$-th covering is given as
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)$ (169)
$=\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})-\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})$ (170)
$\leq\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathrm{Lip}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})-\sup_{h\in\mathcal{N}_{k}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+2\epsilon^{\prime}$ (171)
$\leq\epsilon_{k}+2\epsilon^{\prime},$ (172)
where $f^{\prime}$ and $\epsilon^{\prime}$ is coming from the definition of the supremum, that is, there exists some $f^{\prime}\in\mathcal{F}$ such that $\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\leq\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+\epsilon^{\prime}$ for any $\epsilon^{\prime}\in\real_{+}$. Since we can take $\epsilon^{\prime}$ arbitrarily small, the discretization error is $\epsilon_{k}$.
Next we show that under the $k$-th covering, we can disentangle the complexity of the lipshicit function class by the covering number as follows:
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ (173)
$=\frac{1}{t}\log\exp{t(\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (174)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\exp{t(\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (175)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\exp{t(\sup_{h\in\mathcal{N}_{k}}A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (176)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}\exp{t(A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}$ (177)
$\leq\frac{1}{t}\log\mathbb{E}_{S}\mathbb{E}_{\sigma}\sum_{h\in\mathcal{N}_{k}}\exp{t(A(h)-\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right])}.$ (178)
To evaluate this exponential moment, we use the Mcdiramid’s inequality. The bounded difference coefficient is
$\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})-\left(\sup_{f^{\prime}\in\mathcal{F}}\frac{1}{n}\sum_{i\neq j}^{n}\sigma_{i}\phi(f^{\prime},h,X_{i},Y_{i})+\frac{1}{n}\sigma^{\prime}_{j}\phi(f^{\prime},h,X_{j},Y_{j})\right)$ (179)
$\leq\sup_{f\in\mathcal{F}}\frac{1}{n}(\sigma_{j}\phi(f^{\prime},h,X_{j},Y_{j})-\sigma^{\prime}_{j}\phi(f^{\prime},h,X_{j},Y_{j}))\leq\frac{2}{n}.$ (180)
Thus
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\sup_{h\in\mathcal{N}_{k}}A(h)-\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ $\leq\frac{1}{t}\log\sum_{h\in\mathcal{N}_{k}}\exp{t^{2}n(\frac{2}{n})/8}$ (181)
$=\frac{1}{t}\log|\mathcal{N}_{k}|+\frac{t}{2n},$ (182)
Next, we analyze $\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$. $\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ is the Rademacher complexity of $\phi(f,h,X,Y)$. Moreover from Lemma 7.4 in blasiok2023unify, $\phi(f,h,X,Y)$ is 2-Lipshitz with respect to $h$. Thus by Talagrand’s contraction lemma, $\sup_{h\in\mathcal{N}_{k}}\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})$.
Thus, in conclusion, we have
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{N}_{k}}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\frac{1}{t}\log|\mathcal{N}_{k}|+\frac{t}{2n}.$ (183)
Thus, combining the discretization result
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\sqrt{\frac{2\log|\mathcal{N}_{k}|}{n}}+\epsilon_{k}.$ (184)
Then, following Dudley’s integral approach as shown in Eq. equation 108, we have
$\mathbb{E}_{S}\mathbb{E}_{\sigma}\left[\sup_{h\in\mathrm{Lip}_{L=1}([0,1],[-1,1])}\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\phi(f,h,X_{i},Y_{i})\right]$ (185)
$\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\inf_{\epsilon\geq 0}\left[4\epsilon+12\int_{\epsilon}^{1}\frac{\sqrt{\ln N(\varepsilon^{\prime},\eta_{0},\|\cdot\|_{\infty})}}{\sqrt{n}}d\epsilon^{\prime}\right]$ (186)
$\leq 2\mathfrak{R}_{\mathcal{D},n}(\mathcal{F})+\frac{C}{\sqrt{n}},$ (187)
where we used the covering number estimate of the 1-dimensional Lipschitz functions, see wainwright_2019 for the details.
D.2 Sub-optimal result of the Rademacher complexity
The evaluation of $|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|$ using Rademacher complexity can be carried out in the same manner as in Appendix D.1, except that randomness is now considered over the $2n$ data points from both the test and train sets. Apart from this difference, the argument is identical, and the resulting bound again depends—just as in Appendix D.1—on the complexity of Lipschitz functions through covering arguments, as well as on the complexity of $\mathcal{F}$. This reflects the inherent complexity of the composite function $h(f(X))$ appearing in the definition of smooth CE. Consequently, compared with Theorem 1, where the covering-based proof avoids dependence on the complexity of Lipschitz functions, this approach yields a suboptimal result.
On the other hand, as shown below, it is also possible to derive the Rademacher complexity bound from the covering-number result of Theorem 1. In this case, the dependence on the complexity of Lipschitz functions disappears, but at the cost of introducing an additional $(\log 2n)^{2}$ factor into the Rademacher complexity bound.
Corollary 6.
$\sup_{f\in\mathcal{F}}|\mathrm{smCE}(f,S_{\mathrm{te}})-\mathrm{smCE}(f,S_{\mathrm{tr}})|\leq C_{1}\mathfrak{R}_{\mathcal{D},2n}(\mathcal{F})(\log 2n)^{2}+\frac{C_{2}}{\sqrt{n}}+2\sqrt{\frac{\log\delta^{-1}}{n}}.$ (188)